# Neuromorphic Attitude Estimation and Control

### v1

@@ Abstract
The real-world application of small drones is mostly hampered by energy limitations. Neuromorphic computing promises extremely energy-efficient AI for autonomous flight, but is still challenging to train and deploy on real robots. In order to reap the maximal benefits from neuromorphic computing, it is desired to perform all autonomy functions end-to-end on a single neuromorphic chip, from low-level attitude control to high-level navigation. This research presents the first neuromorphic control system using a spiking neural network (SNN) to effectively map a drone’s raw sensory input directly to motor commands. We apply this method to low-level attitude estimation and control for a quadrotor, deploying the SNN on a tiny Crazyflie. We propose a modular SNN, separately training and then merging estimation and control sub-networks. The SNN is trained with imitation learning, using a flight dataset of sensory-motor pairs. Post-training, the network is deployed on the Crazyflie, issuing control commands from sensor inputs at $500$Hz. Furthermore, for the training procedure we augmented training data by flying a controller with additional excitation and time-shifting the target data to enhance the predictive capabilities of the SNN. On the real drone the perception-to-control SNN tracks attitude commands with an average error of $3$ degrees, compared to $2.5$ degrees for the regular flight stack. We also show the benefits of the proposed learning modifications for reducing the average tracking error and reducing oscillations. Our work shows the feasibility of performing neuromorphic end-to-end control, laying the basis for highly energy-efficient and low-latency neuromorphic autopilots.
@@ I Introduction
A promising solution to these challenges lies in the emerging field of neuromorphic hardware [7]. Neuromorphic systems, including processors and sensors such as event-based cameras [8, 9], draw inspiration from neural systems found in nature. These systems use sparse and asynchronous spikes to transmit information that are both energy-efficient and enable high-speed processing. This approach is particularly well-suited for dynamic environments where rapid decision-making is crucial [10].
@@
Neuromorphic control is a nascent field at the intersection of neuroscience and robotics control theory [13]. The benefits of neuromorphic hardware, such as fast inference and high energy efficiency [14], harmonize with demanding control and estimation tasks. While the output of rudimentary sensors for quadrotors, such as Inertial-Measurement-Units, can already be processed at the high frequencies necessary for agile and robust control, vision-based tasks are severely limited by processing power on a flying machine [15]. However, Dimitrova et al. [16], have shown that using event-based cameras allows a quadrotor to track the horizon at extremely high speeds. To further increase the potential of such a system, the authors of [17] showed that integration of this horizon tracker with a manually-tuned SNN controller on a neuromorphic processor leads to even faster control.
@@
Recent breakthroughs in quadrotor research have achieved impressive results, such as outperforming human pilots in drone races using only onboard computations [19]. Despite these accomplishments, the reliance on slower frame-based vision systems, typically operating at 30Hz or lower, highlights a significant gap where neuromorphic solutions could offer substantial improvements. These examples underscore the critical need for fully integrated neuromorphic systems capable of high-speed data processing.
@@
To allow such a unified system, the entire estimation and control loop needs to be considered. Despite the promising results in partial implementations, a fully integrated neuromorphic system connecting sensor inputs directly to motor commands has not yet been realized in operational quadrotors. Results focusing exclusively on lower-level SNN control have been obtained using manually tuned networks [20, 21] or were limited to simulation [22, 23]. Moreover, even state-of-the-art learned quadrotor controllers using regular Multilayer Perceptrons (MLPs) as presented in [24, 25] and [26], that were learned with Reinforcement Learning (RL), assume full state knowledge or need a lower-level controller to go from rate commands to motor outputs. Zhang et al. [27] have demonstrated in simulation that by using an expert priviliged policy, an MLP can be trained to perform end-to-end control. This privileged information is not always available in the real world, however.
@@
Notable efforts towards a complete end-to-end neuromorphic system include the use of Intel’s Loihi processor [28] in a quadrotor for velocity control based on optical flow estimates from event-based cameras [29], which successfully combined ego-motion estimation with a basic linear controller. The experimental results confirmed the potential of neuromorphic technology, as the vision ran at frequencies between $274$ $1600$Hz, while only spending $7$ mW for network inference. However, this system still required traditional microprocessors for the final control algorithm and lacked a mechanism to correct steady-state errors. In [30] a closed-form spiking network was proposed that could do end-to-end control and estimation for linear systems and was shown to perform well with a small number of neurons in simulation. Since this approach needs to be able to read out a floating point ”firing rate” of neurons in the hidden layer, it is not trivially implemented on commonly available neuromorphic hardware where the input and outputs are limited to vectors of binary spikes.
@@
The main contribution of this article is that we design, train, and implement the first fully neuromorphic control system for attitude estimation and control of quadrotors. The proposed method involves real-time processing from sensors to actuators and does not require traditional computing hardware. Our approach is to train two separate sub-networks, one for state estimation and one for control, and to merge them after training. For both parts of the network, we employ supervised / imitation learning. Especially for the control part, this presents substantial challenges, as the network needs to cope with (i) sensor bias, (ii) delays due to the progressive updates of spiking neural networks, and (iii) the reality gap. Additional contributions of our work concern how we tackled these challenges. For the sensor bias, we find that constraining the parameters of a small group of neurons to function as integrators is necessary for successful training results. For the delays in the SNN, we propose to time-shift the targets for learning, so that the SNN predicts future outputs of the traditional controller. This brings substantial performance improvement. For the reality gap, we first add noise to the motor outputs of the traditional controller to sufficiently excite the system and avoid biases in learning. Subsequently, we gather more training data with a first version of the SNN, so that relevant off-target attitudes and rates are explored. Finally, we evaluate system performance in real-world conditions, comparing the trained SNN with traditional control methods.
@@ II-B1 LIF neurons
where $j$ and $i$ denote presynaptic (input) and postsynaptic (output) neurons within a layer, $k$ the neurons in the same layer as $i$, $s\in\{0,1\}$ a neuron spike, and $w^{ij}$ and $w^{ik}$ feedforward and recurrent connections (if any), respectively. The leak values of the two internal state variables are denoted by $\tau_{i}^{\text{mem}}$ and $\tau_{i}^{\text{syn}}$. A neuron fires an output spike if the membrane potential $\upsilon_{i}$ exceeds threshold $\theta_{i}$ to all connected neurons, resetting its membrane potential to zero at the same time.
@@ II-B2 Combination of networks
To facilitate learning of specific tasks and increase the debugability, the training is split into two parts; estimation and control. By learning layers of spiking neurons that have a certain function, there is more control over the stability of the final solution, and it also reduces the search space. Since we define the input- and output values of both sub-networks as a linear multiplication of the input- or output-vector respectively, the networks can be easily combined. The output of the first network can be written as $y(t)=W_{\text{o}}s(t)$, with $s(t)$ the spikes in the hidden layer, and the input to the next network is $x(t)=W_{\text{i}}y(t)$. We can now combine these by multiplying the weight matrices of the output weights $W_{\text{o}}$ of the first network and the input weights $W_{\text{i}}$ of the second, as introduced in [29], since these are both linear transformations. The attitude-part of the input to the second network can therefore be written as
$\begin{bmatrix}\phi_{\text{est}}\\ \theta_{\text{est}}\\ \psi_{\text{est}}\end{bmatrix}=W_{\text{i}}W_{\text{o}}s(t)$ (3)
Stacking the binary output spikes of the first network with the floating-point command-values that are passed (see Figure 1), the new set of weights to the hidden layer of the second network can be written as:
$W_{\text{new}}=\begin{bmatrix}0&W_{\text{i, command}}\\ W_{\text{i}}W_{\text{o}}&0\\ \end{bmatrix}$ (4)
@@ II-C Training
$\frac{d}{dx}\left(\frac{1}{s}\arctan(sx)\right)=\frac{1}{1+(sx)^{2}}$ (6)
where $s$ is the slope of the surrogate. A higher slope results in a more accurate proxy of the real gradient, but can lead to vanishing gradients for neurons with a very low or high membrane potential. A shallow slope, on the other hand, is less accurate but leads to less ”dead” neurons that have no contribution to the output. Among alternatives for the surrogate-gradient is the derivative of the Sigmoid, but research has shown that the exact shape does not matter [36]. The slope $s$ of the derivative however, does have a large influence on the training speed and final results. For this work, the slope $s$ has been set to 7.
@@ II-C1 Delay in SNN, training with time-shifted data
During training-implementation iterations, oscillations were observed on the real quadrotor. After investigation, these were attributed to a delay in the output of the network versus the target control signal. Due to the nature of the SNN with the implicit memory due to the leaking voltage and current, the output was delayed. This can be observed in Figure 2. In the top part of the figure, the Pearson Correlation between the output of the SNN and the regular PID is compared for different shifts in time on the entire data set. In the bottom part of the figure, a small time sequence is shown that clearly shows the lag. The correlation is highest for 5-6 timesteps shift, indicating that this is indeed a problem when one trains SNNs for highly dynamic tasks that require a quick response to fast changes. In the case of a controller, a small delay in the derivative command will actually induce oscillations.
@@ II-C2 Imitation learning; reducing the reality gap
To address this, we expanded the training data to include a broader, more realistic range of states. Initially, the SNN controller was trained on data generated with the reference controller in the loop, as described in Section II-C. We then conducted additional data collection in two steps to diversify the training set: (1) flying the quadrotor with the initially trained SNN controller in the loop, while simultaneously logging the outputs the reference controller would have provided. This approach exposed the SNN to a set of states it is likely to encounter, fine-tuning the network around these points. (2) Introducing random disturbances to the regular PID controller’s outputs to simulate unexpected environmental or system changes. Specifically, disturbances were applied to pitch, roll, and yaw commands at a $1\%$ probability per timestep (at $500$Hz), lasting 0.2 seconds each, with disturbance size $X\sim U\{0,50\}\%$ of the absolute maximum command.
@@ II-C4 Integrator
To address these issues, we fixed specific parameters of a small subgroup of neurons during training to ensure stability. Specifically, we set the neuron’s leak parameters $\tau_{i}^{\text{syn}}$ and $\tau_{i}^{\text{mem}}$ and threshold $\theta_{i}$ to 1 which enabled it to retain incoming signals without decay, facilitating integration. By training only the input and output weights and averaging spike outputs on integral data alone, we achieved a spike rate approximating the cumulative incoming signal, making the neuron responsive to transient and steady-state inputs. This fixed-parameter integrator provided the needed stability, outperforming the fully trained version and meeting the SNN-based system’s control requirements.
@@ II-D Hardware setup
[…]
@@ III-A Position control
To demonstrate the capabilities of the proposed SNN, we include it in a position control task. The higher level attitude commands together with the IMU values are send as inputs to the SNN, which produces pitch, roll and yaw torque commands. After a short period of hovering at $(x,y)=(0,0)$, the Crazyflie is commanded to move 1 meter in $x$-direction after which it is commanded to move back to $(0,0)$. For both the SNN and PID controller, these tests were performed ten times. In Figure 3, the position control results are shown.
Figure 3: Position step responses of the SNN system (top) and the regular PID flight stack (bottom). The SNN can is able to accurately track the attitude references as given by the outer-loop position controller and maintains a stable flight path.
The results show that performing attitude estimation and control using an onboard SNN results in stable reference tracking, comparable to the regular flight stack of the Crazyflie.
@@ III-B Impact of Time-Shifted and Augmented Training Data on SNN Performance
During testing, it was quickly identified that training the fusion network without augmenting the dataset does not produce a network that is able to be used in flight. Therefore, it was absolutely necessary to augment the dataset for this sub-network. However, to further investigate the behavior of the SNN and the influence of the modifications to the training procedure, another test is performed. Since the directly controlled variable is the attitude command, we compare the response of differently trained networks to an attitude setpoint change. For these tests, the Crazyflie received a roll setpoint of $0\degree$ for 2 seconds, followed by a setpoint of $+10\degree$ for $1.5$ seconds, a setpoint of $-10\degree$ for $1.5$ seconds before returning to a $0\degree$ setpoint for 2.5 seconds. Again, we performed ten tests per controller. The combined results of these ten tests per controller are shown in Figure 4, with A) the final SNN, B) the SNN that was trained on the augmented dataset, C) the SNN that was trained on time-shifted data, but without augmenting the dataset and D) the regular attitude estimator and controller on the Crazyflie.
@@
The Root Mean Square Error (RMSE) between the commanded roll setpoint and the resulting (estimated) roll angle is given in Table I, together with the average standard deviation (SD) of the response with respect to the average of all tests with the same controller. With a tracking error of only $3.03\degree$, the network is able to correctly estimate the attitude and also control it. Adding the suggested modifications to the training procedure reduces the tracking error from $3.24\degree$ to $3.03\degree$ compared to $2.57\degree$ for the reference controller (please note that the reference controller does not have the 2ms delay that is introduced by the UART communication and network inference time). Also, training on time-shifted data significantly reduces the oscillations as can be seen in Figure 4. This can also be inferred from the average SD that is significantly lower for the fully-trained SNN, showing that the controller performs more consistently across multiple tests. On the other hand it is visible that the modifications result in a slightly longer rise-time.
Controller RMSE (deg.) avg. SD
SNN (time-shifted & augmented) 3.03° 0.77
SNN (augmented) 3.10° 0.95
SNN (time-shifted) 3.24° 0.92
SNN (baseline) 3.14° 1.16
PID 2.57° 0.15
TABLE I: Root Mean Square Error (RMSE) and Standard Deviation (SD) comparison between different controllers

### v2

@@ Abstract
The real-world application of small drones is mostly hampered by energy limitations. Neuromorphic computing promises extremely energy-efficient AI for autonomous flight but is still challenging to train and deploy on real robots. To reap the maximal benefits from neuromorphic computing, it is necessary to perform all autonomy functions end-to-end on a single neuromorphic chip, from low-level attitude control to high-level navigation. This research presents the first neuromorphic control system using a spiking neural network (SNN) to effectively map a drone’s raw sensory input directly to motor commands. We apply this method to low-level attitude estimation and control for a quadrotor, deploying the SNN on a tiny Crazyflie. We propose a modular SNN, separately training and then merging estimation and control sub-networks. The SNN is trained with imitation learning, using a flight dataset of sensory-motor pairs. Post-training, the network is deployed on the Crazyflie, issuing control commands from sensor inputs at 500Hz. Furthermore, for the training procedure we augmented training data by flying a controller with additional excitation and time-shifting the target data to enhance the predictive capabilities of the SNN. On the real drone, the perception-to-control SNN tracks attitude commands with an average error of 3.0 degrees, compared to 2.7 degrees for the regular flight stack. We also show the benefits of the proposed learning modifications for reducing the average tracking error and reducing oscillations. Our work shows the feasibility of performing neuromorphic end-to-end control, laying the basis for highly energy-efficient and low-latency neuromorphic autopilots.
Index Terms:
Imitation Learning, Neurorobotics, Machine Learning for Robot Control
@@ I Introduction
A promising solution to these challenges lies in the emerging field of neuromorphic hardware [7]. Neuromorphic systems, including processors and sensors such as event-based cameras [8, 9], draw inspiration from neural systems found in nature. These systems use sparse and asynchronous spikes to transmit information that are both energy-efficient and enable high-speed processing. Due to the low latency, this approach is particularly well-suited for dynamic environments where rapid decision-making is crucial [10].
@@
Neuromorphic control is a nascent field at the intersection of neuroscience and robotics control theory [13]. The benefits of neuromorphic hardware, such as fast inference and high energy efficiency [14], harmonize with demanding control and estimation tasks. While the output of rudimentary sensors for quadrotors, such as Inertial-Measurement-Units, can already be processed at the high frequencies necessary for agile and robust control, vision-based tasks are severely limited by processing power on a flying machine [15]. However, Dimitrova et al. [16] have shown that using event-based cameras allows a quadrotor to track the horizon at extremely high speeds. To further increase the potential of such a system, the authors of [17] showed that integration of this horizon tracker with a manually-tuned SNN controller on a single neuromorphic processor leads to even faster control, benefiting from having all parts on the same chip.
@@
Recent breakthroughs in quadrotor research have achieved impressive results, such as outperforming human pilots in drone races using only onboard computations [19]. Also, Song et al. [4] show that for these tasks, optimal control methods are no longer sufficient and are beaten by Reinforcement Learning (RL) employing Deep Learning.
@@
Despite these accomplishments, the reliance on slower frame-based vision systems, typically operating at 30Hz or lower, highlights a significant gap where neuromorphic solutions could offer substantial improvements. These examples underscore the critical need for fully integrated neuromorphic systems capable of high-speed data processing.
@@
To allow such a unified system, the entire estimation and control loop needs to be considered. Despite the promising results in partial implementations, a fully integrated neuromorphic system connecting sensor inputs directly to motor commands has not yet been realized in operational quadrotors. Results focusing exclusively on lower-level SNN control have been obtained using manually tuned networks [20, 21] or were limited to simulation [22, 23]. Moreover, even state-of-the-art learned quadrotor controllers using regular Multilayer Perceptrons (MLPs) as presented in [24, 25] and [26], that were learned with RL, assume full state knowledge or need a lower-level controller to go from rate commands to motor outputs. Zhang et al. [27] have demonstrated in simulation that by using an expert privileged policy, an MLP can be trained to perform end-to-end control. But also here the observation model, containing the measurements, included a direct measurement of the drone’s attitude. However, such privileged information – complete and accurate state information – is rarely available in real-world scenarios. This limitation is further exacerbated by the reality gap, that arises when algorithms trained or evaluated in simulation must cope with real-world conditions characterized by imperfect measurements, sensor noise, actuator delays, and unpredictable environmental influences.
@@
Notable efforts towards a complete end-to-end neuromorphic system include the use of Intel’s Loihi processor [28] in a quadrotor for velocity control based on optical flow estimates from event-based cameras [29], which successfully combined ego-motion estimation with a basic linear controller. The experimental results confirmed the potential of neuromorphic technology, as the vision ran at frequencies between 274-1600Hz, while only spending 7mW for network inference compared to 14-25Hz on a Jetson Nano that required 1-2W for inference. The neuromorphic system was not only significantly faster, but also required orders of magnitude less power. However, it still relied on a companion computer for attitude control, introducing delays, increasing power consumption, and adding weight to the drone. Moreover, the linear neuromorphic controller lacked a mechanism to compensate for steady-state errors, such as those caused by sensor biases like gyroscope drift. With our work, we want to demonstrate how the pipeline of [29] could be extended to run on a single neuromorphic chip. In [30] a closed-form spiking network was proposed that could do end-to-end control and estimation for linear systems and was shown to perform well with a small number of neurons in simulation. Since this approach needs to be able to read out a floating point ”firing rate” of neurons in the hidden layer, it is not trivially implemented on commonly available neuromorphic hardware where the input and outputs are limited to vectors of binary spikes.
@@
The main contribution of this article is that we design, train, and implement the first fully neuromorphic system for attitude estimation and control of quadrotors. The proposed method involves real-time processing from sensors to actuators and does not require traditional computing hardware. Our approach is to train two separate sub-networks, one for state estimation and one for control, and to merge them after training. For both parts of the network, we employ supervised / imitation learning. In our creation of the training scheme we had to overcome substantial challenges, as the spiking neural network needs to cope with (i) sensor bias, (ii) delays due to the progressive updates of spiking neural networks, (iii) the reality gap and (iv) converting binary spikes to a motor command that leads to smooth control. Additional contributions of our work concern how we tackled these challenges. For the sensor bias, we find that constraining the parameters of a small subgroup of neurons to function as integrators is necessary for successful training results. These integrator neurons can now operate analogously to the integral component of a standard PID controller, effectively mitigating persistent sensor biases. For the delays in the SNN, we propose to time-shift the targets for learning, so that the SNN predicts future outputs of the traditional controller. This brings substantial performance improvement. For the reality gap, we first add noise to the motor outputs of the traditional controller to sufficiently excite the system and avoid biases in learning. Subsequently, we gather more training data with a first version of the SNN, so that relevant off-target attitudes and rates are explored. Finally, we evaluate system performance in real-world conditions, comparing the trained SNN with traditional control methods.
@@ II-B1 LIF neurons
where $j$ and $i$ denote presynaptic (input) and postsynaptic (output) neurons within a layer, $k$ the neurons in the same layer as $i$, $s\in[0,1]$ a neuron spike, and $w^{ij}$ and $w^{ik}$ feedforward and recurrent connections (if any), respectively. The leak values of the two internal state variables are denoted by $\tau_{i}^{\text{mem}}$ and $\tau_{i}^{\text{syn}}$. A neuron fires an output spike if the membrane potential $\upsilon_{i}$ exceeds threshold $\theta_{i}$ to all connected neurons, resetting its membrane potential to zero at the same time.
@@ II-B2 Combination of networks
To facilitate learning of specific tasks and increase the debugability, the training is split into two parts; estimation and control. By learning layers of spiking neurons that have a certain function, there is more control over the stability of the final solution, and it also reduces the search space. Since we define the input- and output values of both sub-networks as a linear multiplication of the input- or output-vector respectively, the networks can be easily combined. The output of the first network can be written as $y(t)=W_{\text{o}}s(t)$, with $s(t)$ the spikes in the hidden layer, and the input to the next network is $x(t)=W_{\text{i}}y(t)$. We can now combine these by multiplying the weight matrices of the output weights $W_{\text{o}}$ of the first network and the input weights $W_{\text{i}}$ of the second, as introduced in [29], since these are both linear transformations. The attitude part of the input to the second network can therefore be written as
$\begin{bmatrix}\phi_{\text{est}}\\ \theta_{\text{est}}\\ \psi_{\text{est}}\end{bmatrix}=W_{\text{i}}W_{\text{o}}s(t).$ (3)
Stacking the binary output spikes of the first network with the floating-point command values that are passed (see Figure 1), the new set of weights to the hidden layer of the second network can be written as
$W_{\text{new}}=\begin{bmatrix}0&W_{\text{i, command}}\\ W_{\text{i}}W_{\text{o}}&0\\ \end{bmatrix}.$ (4)
@@ II-C Training
$\frac{d}{dx}\left(\frac{1}{s}\arctan(sx)\right)=\frac{1}{1+(sx)^{2}},$ (6)
where $s$ is the slope of the surrogate. A higher slope results in a more accurate proxy of the real gradient, but can lead to vanishing gradients for neurons with a very low or high membrane potential. A shallow slope, on the other hand, is less accurate but leads to less ”dead” neurons that have no contribution to the output. Among alternatives for the surrogate gradient is the derivative of the Sigmoid, but research has shown that the exact shape does not matter [36]. The slope $s$ of the derivative, however, does have a large influence on the training speed and final results. For this work, the slope $s$ has been set to 7.
@@ II-C1 Delay in SNN, training with time-shifted data
During training-implementation iterations, oscillations were observed on the real quadrotor. After investigation, these were attributed to a delay in the output of the network versus the target control signal. Due to the nature of the SNN with the implicit memory due to the leaking voltage and current, the output was delayed. This can be observed in Figure 2. In the top part of the figure, the Pearson Correlation between the output of the SNN and the regular PID is compared for different shifts in time on the entire data set. In the bottom part of the figure, a small time sequence is shown that clearly shows the lag. The correlation is highest for 5-6 timesteps shift, indicating that this is indeed a problem when one trains SNNs for highly dynamic tasks that require a quick response to fast changes. In the case of a controller, a small delay in the derivative command will induce oscillations.
@@ II-C2 Imitation learning; reducing the reality gap
To address this, we expanded the training data to include a broader, more realistic range of states. Initially, the SNN controller was trained on data generated with the reference controller in the loop, as described in Section II-C. We then conducted additional data collection in two steps to diversify the training set: (1) flying the quadrotor with the initially trained SNN controller in the loop, while simultaneously logging the outputs the reference controller would have provided. This approach exposed the SNN to a set of states it is likely to encounter, fine-tuning the network around these points. (2) Introducing random disturbances to the regular PID controller’s outputs to simulate unexpected environmental or system changes. Specifically, disturbances were applied to pitch, roll, and yaw commands at a 1% probability per timestep (at 500Hz), lasting 0.2 seconds each, with disturbance size $X\sim U(0,50)\%$ of the absolute maximum command.
@@ II-C4 Integrator
To address these issues for SNNs, we propose fixing certain neuron parameters within a small subgroup of neurons during training to ensure stability. Specifically, we set the leak parameters $\tau_{i}^{\text{syn}}$ and $\tau_{i}^{\text{mem}}$ and threshold $\theta_{i}$ of 10 neurons in the control layer to 1. This allowed the neurons to integrate incoming signals without decay. By training only the input and output weights and averaging spike outputs on integral data alone, we achieved a spike rate approximating the cumulative incoming signal, making the neuron responsive to transient and steady-state inputs. This approach is validated in Figure 3, which compares training curves for an integration task with fixed versus free neuron leak and threshold parameters. The fixed-parameter integrator provided the necessary stability, outperforming the fully unconstrained trained approach and satisfying the SNN-based system’s control requirements.
Figure 3: Training loss curves comparing fixed versus free neuron leak and threshold parameters. The proposed approach of fixing neuron parameters leads to stable convergence during training. Allowing these parameters to remain free results in training becoming trapped in local minima.
@@ II-D Hardware setup
To demonstrate the capabilities of our approach, we have implemented it in the control loop of the tiny open-source quadrotor Crazyflie [32]. By adding a Teensy 4.0 development board to the Crazyflie, the necessary computation power for running an SNN on a processor was obtained. This allowed us to run the complete SNN from input encoding to control commands at 500Hz in C++ on the ARM Cortex-M7 microprocessor. To carry the extra weight of the Teensy, the regular 16mm brushed motors of the Crazyflie are swapped with 20mm brushed motors. To maximize the accuracy of the network while utilizing the Teensy to its full extent, the network was optimized for speed by removing unnecessary neurons. This was done by performing inference on a number of test sequences and calculating the total contribution of a neuron on the output by calculating the total number of spikes emitted multiplied by its weight to all outputs. Now the $N$ lowest contributing neurons can be removed from the implementation in C++ on the Teensy. Although the network was trained with 150-150-130 neurons per layer respectively, we reduced the size to 150, 100, and 80 per layer respectively. By mainly pruning the neurons with recurrent connections this way, we almost halve the number of mathematical operations while retaining over 99% of the original MSE that was used as the loss function during training.
@@ III-A Position control
To demonstrate the capabilities of the proposed SNN, we include it in a position control task. The higher-level attitude commands together with the IMU values are sent as inputs to the SNN, which produces pitch, roll and yaw torque commands. After a short period of hovering at $(x,y)=(0,0)$, the Crazyflie is commanded to move 1 meter in $x$-direction after which it is commanded to move back to $(0,0)$. For both the SNN and PID controller, these tests were performed ten times. In Figure 4, the position control results are shown. The results show that performing attitude estimation and control using an onboard SNN results in stable reference tracking, comparable to the regular flight stack of the Crazyflie.
Figure 4: Position step responses of the SNN system (top) and the regular PID flight stack (bottom) for 10 individual test runs. The SNN can accurately track the attitude references as given by the outer-loop position controller and maintain a stable flight path.
@@ III-B Impact of Time-Shifted and Augmented Training Data on SNN Performance
During testing, it was quickly identified that training the fusion network without augmenting the dataset does not produce a network that can be used in flight. Therefore, it was necessary to augment the dataset for this sub-network. However, to further investigate the behavior of the SNN and the influence of the modifications to the training procedure, another test is performed. Since the directly controlled variable is the attitude command, we compare the response of differently trained networks to an attitude setpoint change. For these tests, the Crazyflie received a roll setpoint of $0\degree$ for 2 seconds, followed by a setpoint of $+10\degree$ for 1.5 seconds, a setpoint of $-10\degree$ for 1.5 seconds before returning to a $0\degree$ setpoint for 2.5 seconds. Again, we performed ten tests per controller. The combined results of these ten tests per controller are shown in Figure 5, with A) the final SNN, B) the SNN that was trained on the augmented dataset, C) the SNN that was trained on time-shifted data, but without augmenting the dataset and D) the regular attitude estimator and controller on the Crazyflie. The Root Mean Square Error (RMSE) between the commanded roll setpoint and the resulting (estimated) roll angle is given in Table I, together with the average standard deviation (SD) of the response with respect to the average of all tests with the same controller. With a tracking error of only $3.03\degree$, the network is able to correctly estimate the attitude and also control it. Adding the suggested modifications to the training procedure reduces the tracking error from $3.24\degree$ to $3.03\degree$ compared to $2.67\degree$ for the reference controller (please note that the reference controller receives the estimated attitude directly, while the SNN needs to internally calculate this). Also, training on time-shifted data significantly reduces the oscillations as can be seen in Figure 5. This can also be inferred from the average SD that is significantly lower for the fully-trained SNN, showing that the controller performs more consistently across multiple tests. On the other hand, training on time-shifted data very slightly increases the rise-time (see Table I). Since the increase is in the order of milli-seconds, it will not affect tasks like obstacle avoidance that generally operate in the 20-40Hz range [41] but it should be considered if it is used in super agile flight.
Controller RMSE avg. SD avg. RT
SNN (time-shifted & augm.) 3.03° 0.77 145ms
SNN (augmented) 3.10° 0.95 130ms
SNN (time-shifted) 3.24° 0.92 145ms
SNN (baseline) 3.14° 1.16 135ms
PID 2.67° 0.23 125ms
TABLE I: Root Mean Square Error (RMSE), Standard Deviation (SD) and rise-time (RT) comparison between different controllers. Note that the PID receives the estimated attitude as input, while the SNN needs to calculate this internally.